OpenStudy (anonymous):

each individual letter of the word TENNESSEE is placed on a piece of paper and all 9 pieces of paper are place in a hat. If one letter is selected at random from the hat, determine the probability that . problem the letter T or N is selected.

OpenStudy (anonymous):

probability of getting a T. there is 1 T and 9 letters so 1/9 probability of getting an N. there's 2 N's and 9 letters so 2/9 probability of getting either (just add probability of each) (1/9)+(2/9)=(1+2)/9=3/9=1/3